OpenStudy (anonymous):

Find the value of x when x+3/9 = x-1/x-3

OpenStudy (anonymous):

satellite please help me

OpenStudy (anonymous):

\[(x+3/9)*(x-3)=x-1\] \[x^2-24/9x-1=x-1\] \[x^2-2.667x-1=x-1\] \[x^2-3.667x=0\] \[x(x-3.667)=0\] x(1)=0 x(2)=3.667

OpenStudy (anonymous):